OpenStudy (anonymous):

Solving an equation involving logarithms: \[(\log_3 x)^2 = \log_3 x^2 + 3\] Specifically, I am looking on guidance on what to do with the left hand side. I'm not sure if there are any rules that govern multiplying logs, though I don't think there are.

OpenStudy (turingtest):

\[(\log_3x)^2=\log_3(x^2)+3\]is this right?

OpenStudy (anonymous):

Yep.

OpenStudy (turingtest):

let\[u=\log_3x\]then we have\[u^2=2u+3\]which is a quadratic we can solve for u from that we can get x

OpenStudy (anonymous):

\[\log_{3}x^{2} = 2 \log_{x}3 \] Therefor: \[(log_{3}x)^{2} \ = 2 \log_{x}3 + 3\]\[(log_{3}x)^{2} \ - 2 \log_{x}3 - 3=0\]Assume: \[(log_{3}x) \ = a\]So: a^2 -2a - 3 = 0 (a-3)(a+1)=0 So logx = 3 and logx = -1

OpenStudy (anonymous):

Thanks to both of you! That gave me the correct answers of 1/3 and 27.